OpenStudy (anonymous):

if y varies directly as x and y=2/3 when x=1/2, find an equation of direct variation the choices are: a. y=3/4x b. y=-4/3x c. y=4/3x d. y=-3/4x

OpenStudy (nikato):

So y=kx And you know y=2/3 and x=1/2 So 2/3= k(1/2) Can you find k?

OpenStudy (anonymous):

no >.<

OpenStudy (raden):

well i just continue from nikato. 2/3 = k * 1/2 times 2 on both sides, get 2/3 * 2 = k * 1/2 * 2 4/3 = k now plug in that k = 4/3 into equation of y=kx

OpenStudy (raden):